OpenStudy (anonymous):

5!6!8! what the answer the question!

OpenStudy (kropot72):

Does your calculator do factorials?

OpenStudy (kropot72):

5! = 5 *4 *3 * 2 * 1 = 120

OpenStudy (kropot72):

6! = 6 * 5 *4 * 3 * 2 * 1 = 720

OpenStudy (kropot72):

8! = 8 * 7 * 6 * 5 *4 * 3 * 2 * 1 = 40320 Now you can easily multiply 5! * 6! * 8!

OpenStudy (kropot72):

@RISKI Are you there?

OpenStudy (unklerhaukus):

\[5!6!8!\\=5!\times6\times5!\times8\times7\times6\times5!\\=8\times7\times6^2\times5!^3\]